# Continuity and equivariant dimension

### v1

@@ 3 Local-Triviality Dimension and Fields
is upper semicontinuous, so there exists a neighborhood $U$ of $x$ such that for each $y\in U$, both $\|ab-1\|_{y}<1$ and $\|ba-1\|_{y}<1$. In particular, $\pi_{y}(a)$ is invertible for all $y\in U$. Since $\sum\limits_{i=0}^{n}\overline{\rho}_{i}(t)$ is invertible modulo $J_{y}$ for all $y\in U$, this shows that
@@
finishing the proof. normal-$\blacksquare$
@@
where ${\mathbb{S}}^{1}$ is a maximal torus: see [29, Theorems 1 and 2] and [28, Theorem III.1.11.1], for $G=SU(2)$, as recalled in [50, Introduction].
@@
    resulting from realizing ${\mathbb{C}}^{n}$ as the irreducible $n$-dimensional $SU(2)$-representation (see [53, Theorem VIII.4.2]).
@@
The $A_{\theta}$ are the fibers of a continuous field by [49, Theorem 8.13 and Example 10.2], and all admit ergodic actions (see [20, p.537]) by the ordinary 2-torus ${\mathbb{T}}^{2}\cong({\mathbb{S}}^{1})^{2}$. This constitutes a continuous field of actions in the sense of [46, Definition 3.1]. Note the following.
@@
  • On the other hand, for non-integral values of $\theta$, none of the dimensions can vanish. The ${\mathbb{T}}^{2}$-action is ergodic, so the eigenspaces for the characters of ${\mathbb{T}}^{2}$ are all 1-dimensional (see [58, Theorem 1]). Therefore an equivariant morphism
@@
Among 3-7, the only nontrivial implication is c normal-$\Longrightarrow$ a. Consider morphisms
@@
Recall from [59, Corollary 4.1] that $\varphi$ uniquely retrieves a completely positive contractive order-zero [59, Definition 2.3] map
@@
Note that $\widetilde{\varphi}$ is ${\mathbb{G}}$-equivariant in the usual sense, and by [59, Theorem 3.3] it is a scaling of a morphism $C({\mathbb{G}})\to A$ by the ${\mathbb{G}}$-invariant, positive, non-zero element $h:=\widetilde{\varphi}(1)\in A$. Since ergodicity of the action implies that $h$ is a scalar, we can divide by it to recover a ${\mathbb{G}}$-equivariant $C^{*}$-morphism $C({\mathbb{G}})\to A$.
Observe also that a ${\mathbb{G}}$-equivariant $C^{*}$-morphism is by necessity one-to-one on the dense Hopf ***-algebra [36, §1.6] ${\mathcal{O}}({\mathbb{G}})\leq C({\mathbb{G}})$: the ${\mathbb{G}}$-invariant state
{tikzpicture}$\begin{tikzpicture}\quad\psi:=\text{state on }A,\quad h_{{\mathbb{G}}}:=\text{the Haar state on ${\mathbb{G}}$}$
restricts to $h_{{\mathbb{G}}}$ along the assumed morphism $C({\mathbb{G}})\to A$, and $h_{{\mathbb{G}}}$ is faithful on ${\mathcal{O}}({\mathbb{G}})$ [36, Corollary 1.7.5]. If the action is 0-dimensional, then, the multiplicity of a ${\mathbb{G}}$-representation $\rho$ in $A$ is precisely $\dim\rho$. For ergodic actions of Kac groups this is also an upper bound for that multiplicity [7, Theorem 17] (see also [58, Theorem 1] for classical compact groups), hence the implication a normal-$\Rightarrow$ d for Kac-type ${\mathbb{G}}$. normal-$\blacksquare$
@@
by [51, Theorem 9.74 and Theorem following Example 9.65]. Next, [33, Theorem 8.1.7] once more shows that the induced grading by the Pontryagin-dual group $\Gamma$ is strong:
@@ 4 Rational tori and spheres
For $k=2n-1$, each generator $z_{j}=x_{j}+iy_{j}$ of $C({\mathbb{S}}^{2n-1}_{\theta})$ is normal and odd, so the identity $1=\sum\limits_{j=1}^{n}z_{j}z_{j}^{*}=\sum\limits_{j=1}^{n}x_{j}^{2}+y_{j}^{2}$ shows that the weak and plain local-triviality dimensions are at most $12n-$ 1. A similar argument applies to $k=2n-2$, with the final generator $z_{n}=x_{n}$ self-adjoint. normal-$\blacksquare$
@@
Now, assume that each $\theta_{jk}$ is a rational number with odd denominator, and suppose $k=2n-1$ is odd. There exists an odd positive integer $M$ such that in the associated quantum torus $C(\mathbb{T}^{n}_{\theta})$, the elements $u_{1}^{M},\ldots,u_{n}^{M}$ are central. The functions $f_{i}:S^{n-1}_{+}\to C(\mathbb{T}^{n}_{\theta})$ defined by
@@
If $k$ is even, then the final generator is self-adjoint, so a similar argument applies with one fewer term. normal-$\blacksquare$
@@
written in lowest terms. From [45, Theorem 3.12] (or [20, Proposition 12.2]) we see that, in this case, there is a rank-$q$ vector bundle ${\mathcal{E}}$ on ${\mathbb{T}}^{2}$ such that
@@
    is identified in [45, Proposition 3.8 and Theorem 3.9] with an integer [^1] The bundle ${\mathcal{E}}$ is only unique up to tensoring with a line bundle [15, Théorème 9], hence the uniqueness of $a$ only modulo $q$. This is also compatible with the fact that $A_{\theta}$ only depends on the coset of $\theta$ modulo ${\mathbb{Z}}$. $a$ for which $ap+bq=1$ (the twist of [45, p.299] is $-a$). In particular, that Chern class cannot vanish.
    Additionally, vector bundles over compact connected Riemann surfaces are completely determined by their ranks and degrees (i.e. Chern classes $c_{1}$) by [56, Proposition, p.2]. Flat bundles on ${\mathbb{T}}^{2}$, for that reason, are topologically trivial. They need not be holomorphically trivial, but ‘topologically’ is what matters in the present context.
@@
    What does hold, and what was perhaps meant by the flatness claim in the paragraph preceding [26, Exercise 2.8], is that the bundle
    ${\mathcal{E}}nd({\mathcal{E}})\cong{\mathcal{E}}\otimes{\mathcal{E}}^{*}$
    itself is flat, and in fact it is topologically trivial as a vector bundle (but not as a bundle of algebras!). The first Chern class $c_{1}({\mathcal{E}}\otimes{\mathcal{E}}^{*})\in H^{2}({\mathbb{T}}^{2},{\mathbb{Z}})$ vanishes by [27, (1.14)], whence triviality by [56, Proposition, p.2] again.
@@
Let $\theta=\frac{p}{q}$ be a lowest-terms rational number.
@@
by 4-3. On the other hand, evaluating the $f$ of 4-6 at either boundary point produces a function algebra $C({\mathbb{S}}^{1})$ generated, respectively, by one of the unitaries $u,v\in C({\mathbb{T}}^{2}_{\theta})$. Focusing on $u$ to fix the notation, further evaluation at a point $z$ of the corresponding circle ${\mathbb{S}}^{1}$ means imposing the relation $u^{q}=z$. The $C^{*}$-algebra generated by a unitary $u$ subject to the relation $u^{q}=z\in{\mathbb{S}}^{1}$ is indeed abelian and $q$-dimensional, and we are done. normal-$\blacksquare$
@@
The space of ordered splittings of ${\mathbb{C}}^{q}$ as a direct sum of lines is the full flag manifold [2, §3] $U(q)/U(1)^{q}$, easily proven simply-connected via the long exact homotopy sequence [54, §17.3]. The quotient by the cyclic ${\mathbb{Z}}/q$-action then produces a space ${\mathbb{P}}_{q}$ with fundamental group ${\mathbb{Z}}/q$, and hence a ${\mathbb{P}}_{q}$-bundle ${\mathcal{E}}_{q}$ over ${\mathbb{T}}^{2}\times(0,1)$. Moreover, the $x_{j}$ jointly give, as described above, a single section of ${\mathcal{E}}_{q}$ over all of ${\mathbb{T}}^{2}\times(0,1)$.
@@
The previous claim results in a contradiction. If one restricts everything in sight to the circle ${\mathbb{S}}^{1}\subset{\mathbb{T}}^{2}$ corresponding to $u$, sections of ${\mathcal{E}}_{q}$ are nothing but loops in the fiber ${\mathbb{P}}_{q}$ over $1\in{\mathbb{S}}^{1}$. As such, $s_{u}$ is not nullhomotopic (as it generates $\pi_{1}({\mathbb{P}}_{q})\cong{\mathbb{Z}}/q$), but $s_{v}$ is. normal-$\blacksquare$
@@
[…]
@@
[…]
@@
The same Example 3.18 also argues that for $S_{\theta}$ the dimensions cannot vanish; the strongest claim that $\mathrm{dim}^{{\mathbb{Z}}/2}_{\mathrm{WLT}}(A_{\theta})\neq 0$ follows from [39, Proposition 1.9], hence the conclusion. normal-$\blacksquare$

### v2

@@ 3 Local-Triviality Dimension and Fields
is upper semicontinuous, so there exists a neighborhood $U$ of $x$ such that for each $y\in U$, both $\|ab-1\|_{y}<1$ and $\|ba-1\|_{y}<1$. In particular, $\pi_{y}(a)$ is invertible for all $y\in U$. Since $\sum\limits_{i=0}^{n}\overline{\rho}_{i}(t\otimes 1)$ is invertible modulo $J_{y}$ for all $y\in U$, this shows that
@@
finishing the proof. $\blacksquare$
@@
where ${\mathbb{S}}^{1}$ is a maximal torus: see [29, Theorems 1 and 2] and [28, Theorem III.1.11.1], for $G=SU(2)$, as recalled in [51, Introduction].
@@
    resulting from realizing ${\mathbb{C}}^{n}$ as the irreducible $n$-dimensional $SU(2)$-representation (see [54, Theorem VIII.4.2]).
@@
The $A_{\theta}$ are the fibers of a continuous field by [50, Theorem 8.13 and Example 10.2], and all admit ergodic actions (see [20, p.537]) by the ordinary 2-torus ${\mathbb{T}}^{2}\cong({\mathbb{S}}^{1})^{2}$. This constitutes a continuous field of actions in the sense of [47, Definition 3.1]. Note the following.
@@
  • On the other hand, for non-integral values of $\theta$, none of the dimensions can vanish. The ${\mathbb{T}}^{2}$-action is ergodic, so the eigenspaces for the characters of ${\mathbb{T}}^{2}$ are all 1-dimensional (see [59, Theorem 1]). Therefore an equivariant morphism
@@
Among 3-7, the only nontrivial implication is c $\Longrightarrow$ a. Consider morphisms
@@
Recall from [60, Corollary 4.1] that $\varphi$ uniquely retrieves a completely positive contractive order-zero [60, Definition 2.3] map
@@
Note that $\widetilde{\varphi}$ is ${\mathbb{G}}$-equivariant in the usual sense, and by [60, Theorem 3.3] it is a scaling of a morphism $C({\mathbb{G}})\to A$ by the ${\mathbb{G}}$-invariant, positive, non-zero element $h:=\widetilde{\varphi}(1)\in A$. Since ergodicity of the action implies that $h$ is a scalar, we can divide by it to recover a ${\mathbb{G}}$-equivariant $C^{*}$-morphism $C({\mathbb{G}})\to A$.
Observe also that a ${\mathbb{G}}$-equivariant $C^{*}$-morphism is by necessity one-to-one on the dense Hopf *-algebra [36, §1.6] ${\mathcal{O}}({\mathbb{G}})\leq C({\mathbb{G}})$: the ${\mathbb{G}}$-invariant state
[…]
restricts to $h_{{\mathbb{G}}}$ along the assumed morphism $C({\mathbb{G}})\to A$, and $h_{{\mathbb{G}}}$ is faithful on ${\mathcal{O}}({\mathbb{G}})$ [36, Corollary 1.7.5]. If the action is 0-dimensional, then, the multiplicity of a ${\mathbb{G}}$-representation $\rho$ in $A$ is precisely $\dim\rho$. For ergodic actions of Kac groups this is also an upper bound for that multiplicity [7, Theorem 17] (see also [59, Theorem 1] for classical compact groups), hence the implication a $\Rightarrow$ d for Kac-type ${\mathbb{G}}$. $\blacksquare$
@@
by [52, Theorem 9.74 and Theorem following Example 9.65]. Next, [33, Theorem 8.1.7] once more shows that the induced grading by the Pontryagin-dual group $\Gamma$ is strong:
@@ 4 Rational tori and spheres
For $k=2n-1$, each generator $z_{j}=x_{j}+iy_{j}$ of $C({\mathbb{S}}^{2n-1}_{\theta})$ is normal and odd, so the identity $1=\sum\limits_{j=1}^{n}z_{j}z_{j}^{*}=\sum\limits_{j=1}^{n}x_{j}^{2}+y_{j}^{2}$ shows that the weak and plain local-triviality dimensions are at most $12n-$ 1. A similar argument applies to $k=2n-2$, with the final generator $z_{n}=x_{n}$ self-adjoint. $\blacksquare$
@@
Now, assume that each $\theta_{jk}$ is a rational number with odd denominator, and suppose $k=2n-1$ is odd. There exists an odd positive integer $M$ such that in the associated quantum torus $C(\mathbb{T}^{n}_{\theta})$, the elements $u_{1}^{M},\ldots,u_{n}^{M}$ are central. The functions $f_{i}:{\mathbb{S}}^{n-1}_{+}\to C(\mathbb{T}^{n}_{\theta})$ defined by
@@
If $k$ is even, then the final generator is self-adjoint, so a similar argument applies with one fewer term. $\blacksquare$
@@
written in lowest terms. From [46, Theorem 3.12] (or [20, Proposition 12.2]) we see that, in this case, there is a rank-$q$ vector bundle ${\mathcal{E}}$ on ${\mathbb{T}}^{2}$ such that
@@
    is identified in [46, Proposition 3.8 and Theorem 3.9] with an integer $a$ for which $ap+bq=1$ (the twist of [46, p.299] is $-a$). In particular, that Chern class cannot vanish. Note that the vector bundle ${\mathcal{E}}$ is only unique up to the tensor product with a line bundle [15, Théorème 9], so $a$ is only unique modulo $q$. This is also compatible with the fact that $A_{\theta}$ only depends on the coset of $\theta$ modulo ${\mathbb{Z}}$.
    Additionally, vector bundles over compact connected Riemann surfaces are completely determined by their ranks and degrees (i.e. Chern classes $c_{1}$) by [57, Proposition, p.2]. Flat bundles on ${\mathbb{T}}^{2}$, for that reason, are topologically trivial. They need not be holomorphically trivial, but ‘topologically’ is what matters in the present context.
@@
    What does hold, and what was perhaps meant by the flatness claim in the paragraph preceding [26, Exercise 2.8], is that the endomorphism bundle ${\mathcal{E}}\otimes{\mathcal{E}}^{*}$ (recovering $A_{\theta}$ as the section algebra 4-3) is flat, and in fact is topologically trivial as a vector bundle (but not as a bundle of algebras!). The first Chern class $c_{1}({\mathcal{E}}\otimes{\mathcal{E}}^{*})\in H^{2}({\mathbb{T}}^{2},{\mathbb{Z}})$ vanishes by [27, (1.14)], whence triviality by [57, Proposition, p.2] again.
@@
Let $\theta=\frac{p}{q}$ be a lowest-terms rational number. Then the following claims hold.
@@
by 4-3. On the other hand, evaluating the $f$ of 4-6 at either boundary point produces a function algebra $C({\mathbb{S}}^{1})$ generated, respectively, by one of the unitaries $u,v\in C({\mathbb{T}}^{2}_{\theta})$. Focusing on $u$ to fix the notation, further evaluation at a point $z$ of the corresponding circle ${\mathbb{S}}^{1}$ means imposing the relation $u^{q}=z$. The $C^{*}$-algebra generated by a unitary $u$ subject to the relation $u^{q}=z\in{\mathbb{S}}^{1}$ is indeed abelian and $q$-dimensional, and we are done. $\blacksquare$
@@
The space of ordered splittings of ${\mathbb{C}}^{q}$ as a direct sum of lines is the full flag manifold [2, §3] $U(q)/U(1)^{q}$, easily proven simply-connected via the long exact homotopy sequence [55, §17.3]. The quotient by the cyclic ${\mathbb{Z}}/q$-action then produces a space ${\mathbb{P}}_{q}$ with fundamental group ${\mathbb{Z}}/q$, and hence a ${\mathbb{P}}_{q}$-bundle ${\mathcal{E}}_{q}$ over ${\mathbb{T}}^{2}\times(0,1)$. Moreover, the $x_{j}$ jointly give, as described above, a single section of ${\mathcal{E}}_{q}$ over all of ${\mathbb{T}}^{2}\times(0,1)$.
@@
The previous claim results in a contradiction. If one restricts everything in sight to the circle ${\mathbb{S}}^{1}\subset{\mathbb{T}}^{2}$ corresponding to $u$, sections of ${\mathcal{E}}_{q}$ are nothing but loops in the fiber ${\mathbb{P}}_{q}$ over $1\in{\mathbb{S}}^{1}$. As such, $s_{u}$ is not nullhomotopic (as it generates $\pi_{1}({\mathbb{P}}_{q})\cong{\mathbb{Z}}/q$), but $s_{v}$ is. $\blacksquare$
@@
$\mathrm{dim}^{{\mathbb{Z}}/2}_{\mathrm{WLT}}(A_{\theta})$ $=\mathrm{dim}^{{\mathbb{Z}}/2}_{\mathrm{LT}}(A_{\theta})=1\quad\text{and}$ (4-9)
@@
$\mathrm{dim}^{{\mathbb{Z}}/2}_{\mathrm{WLT}}(A^{\theta})\leq\mathrm{dim}^{{\mathbb{Z}}/2}_{\mathrm{WLT}}(S^{\theta})\quad\text{and}\quad\mathrm{dim}^{{\mathbb{Z}}/2}_{\mathrm{LT}}(A^{\theta})\leq\mathrm{dim}^{{\mathbb{Z}}/2}_{\mathrm{LT}}(S^{\theta}).$
@@
The same Example 3.18 also argues that for $S_{\theta}$ the dimensions cannot vanish; the strongest claim that $\mathrm{dim}^{{\mathbb{Z}}/2}_{\mathrm{WLT}}(A_{\theta})\neq 0$ follows from [39, Proposition 1.9], hence the conclusion. $\blacksquare$
@@
Acknowledgments
The views expressed in this document are those of the authors and do not reflect the official policy or position of the U.S. Department of Defense or the U.S. Government.
The authors would like to thank the referee for helpful comments.
A.C. was partially supported by NSF grant DMS-2001128.
This work is part of the project Graph Algebras partially supported by EU grant HORIZON-MSCA-SE-2021 Project 101086394.